OpenStudy (anonymous):

A swimming pool can be filled in 19 hours if water enters through a pipe alone, or in 20 hours if water enters through a hose alone. If water is entering through both the pipe and the hose, how long will it take to fill the pool ?

OpenStudy (anonymous):

I know a great strategy

OpenStudy (anonymous):

first u would find how much wan done in 1 hour

OpenStudy (anonymous):

\[pipe=\frac{ pool }{ 19 h }\]\[hose=\frac{ pool }{ 20 h }\] \[pipe+hose=\frac{ pool }{ 19 h }+\frac{ pool }{ 20 h }=\frac{ pool \times 20 }{ 19 h \times 20}+\frac{ pool \times 19}{ 20 h \times 19 }\] \[=\frac{ pool \times 20 + pool \times 19}{ 380 h} =\frac{39pool}{380h}=\frac{pool}{9.74h}.\]Takes 39/380 = 9.75 hours to fill pool using both at the same time.

OpenStudy (anonymous):

next take the reciprocal which is 380/39

OpenStudy (anonymous):

this is about 9.75